OpenStudy (anonymous):

solve for x 6(x-4)-6=-3(-6x+4)-2x

OpenStudy (anonymous):

explain as well please, i get x = 18 not sure if correct

OpenStudy (anonymous):

6x - 24 - 6 = 18x - 12 - 2x 6x - 30 = 16x - 12 10x = -18 x = -1.8

OpenStudy (anonymous):

I agree with moneybird's answer

OpenStudy (mertsj):

That's because moneybird is correct.